# HyperLoader: Integrating Hypernetwork-Based LoRA and Adapter Layers into Multi-Task Transformers for Sequence Labelling

### v2

@@ 6 Conclusions
We proposed a method that combines parameter-efficient methods in a multi-task setting. Using a hypernetwork, we generate the weights of the adapter layers and LoRA matrices conditioned on the task, the transformer layer of the model, and the position of the adapter and LoRA matrices within this layer. Our method outperforms previous work in single-task and multi-task fine-tuning that combines different parameter-efficient methods and hypernetwork-based approaches using the full training and validation splits of each dataset and in a low-resource configuration using only 10% and 20% of those partitions to train the model. We provide empirical evidence that the improvement in performance is not only because of augmenting the trainable parameters since the hyperdecoder model has the largest number of trainable weights. Therefore, the combination of parameter-efficient methods and weight generation is a strong alternative to solving Sequence Labelling tasks in a multi-task setting.
@@ 7 Limitations and Future work
Although our model performs strongly compared to previous work in multi-task settings using a parameter-efficient fine-tuning approach that reduces negative transfer, underfitting, and overfitting across tasks, it still needs to have access to all datasets during training and requires complete re-training if a new task is added. Curriculum learning Bengio et al. (2009); Wang et al. (2021); Soviany et al. (2022); Piergiovanni et al. (2023) is a promising direction for future work and could alleviate the limitation above. This technique improves the model’s learning efficiency, often leading to faster convergence and better final performance.

### v3

@@ 6 Conclusions
We proposed a method that combines parameter-efficient methods in a multi-task setting. Using a hypernetwork, we generate the weights of the adapter layers and LoRA matrices conditioned on the task, the transformer layer of the model, and the position of these modules within this layer. Our method outperforms previous work in single-task and multi-task fine-tuning that combines different parameter-efficient methods and hypernetwork-based approaches using the full training and validation splits of each dataset and in a low-resource configuration using only 10% and 20% of those partitions to train the model. We provide empirical evidence that the improvement in performance is not only because of augmenting the trainable parameters since the hyperdecoder model has the largest number of trainable weights. Therefore, the combination of parameter-efficient methods and weight generation is a strong alternative to solving Sequence Labelling tasks in a multi-task setting.
@@ 7 Limitations and Future work
Our model excels in multi-task settings with a parameter-efficient fine-tuning approach that mitigates negative transfer, underfitting, and overfitting. However, it still requires access to all datasets during training and needs complete retraining when a new task is added. Curriculum learning Bengio et al. (2009); Wang et al. (2021); Soviany et al. (2022); Piergiovanni et al. (2023) could address this limitation by enhancing learning efficiency, potentially leading to faster convergence and improved performance.
Appendix A Corpora labels
Table 5 shows each Sequence Labelling dataset’s complete list of labels used to train and evaluate the HyperLoader model. Based on the labels’ names, we calculate label overlap between the datasets. Nine different labels are shared across the datasets, representing 3.78% of the total labels.
ATIS labels
fromloc.airport_name arrive_time.start_time flight_number cost_relative
connect flight_days restriction_code depart_date.date_relative
return_date.month_name mod arrive_date.month_name city_name
depart_date.day_number compartment depart_time.start_time airline_name
meal depart_date.month_name time_relative return_date.today_relative
depart_time.period_mod flight_mod airport_name stoploc.airport_code
depart_date.year fare_basis_code today_relative airport_code
fromloc.state_name toloc.city_name economy booking_class
arrive_date.today_relative arrive_date.date_relative toloc.airport_code fromloc.airport_code
day_number stoploc.city_name state_code month_name
arrive_date.day_name arrive_time.period_of_day state_name aircraft_code
period_of_day return_time.period_mod day_name stoploc.state_code
toloc.state_code depart_time.time_relative toloc.airport_name return_date.date_relative
fromloc.city_name return_date.day_number depart_time.time depart_date.day_name
arrive_time.time meal_code or class_type
return_date.day_name time toloc.state_name arrive_date.day_number
days_code arrive_time.period_mod arrive_time.time_relative flight_stop
depart_time.period_of_day transport_type round_trip meal_description
fare_amount toloc.country_name arrive_time.end_time depart_time.end_time
flight fromloc.state_code depart_date.today_relative flight_time
airline_code return_time.period_of_day stoploc.airport_name
mTOP labels
school contact_related recipes_diet news_source
todo recipes_unit_nutrition person_reminded attendee
recipes_source date_time news_topic music_album_title
life_event contact_method recipes_time_preparation type_contact
news_reference similarity name_app recipes_qualifier_nutrition
recipes_cooking_method timer_name contact_removed employer
recipes_excluded_ingredient method_recipes type_relation group
news_type content_exact ordinal news_category
recipes_unit_measurement user_attendee_event music_playlist_title sender
music_provider_name recipes_dish location amount
music_rewind_time music_type alarm_name weather_temperature_unit
music_album_modifier category_event education_degree recipes_meal
period music_artist_name music_radio_id method_timer
recipes_cuisine phone_number music_genre weather_attribute
music_track_title recipes_included_ingredient attribute_event method_retrieval_reminder
major contact_added recipes_rating contact
gender age recipient job
recipes_type recipes_type_nutrition attendee_event recipes_attribute
music_playlist_modifier title_event type_content
SNIPS labels
object_location_type poi state genre
object_name album playlist movie_type
movie_name artist restaurant_name restaurant_type
spatial_relation timerange object_part_of_series_type current_location
served_dish city object_select music_item
country cuisine geographic_poi sort
condition_temperature object_type party_size_description service
track entity_name party_size_number rating_value
playlist_owner condition_description rating_unit location_name
year facility best_rating
Movie labels
genre rating year plot
ratings_average director song title
trailer review actor character
MovieTrivia labels
award relationship quote genre
character_name director plot year
soundtrack origin actor opinion
mTOD labels
demonstrative_reference datetime weather negation
alarm news timer reminder
location
Restaurant labels
restaurant_name cuisine rating price
dish hours amenity location
Table 5: List of labels for each used Sequence labelling dataset to evaluate our proposed approach.